# Calibration-free single-frame super-resolution fluorescence microscopy

### v2

@@ IV.1 Experiment
The PVA forms an environment where terrylene exhibits very limited photostability, with some of the individual molecules undergoing several blinking cycles before irreversibly photobleaching. We captured a sequence of images from the same location within the sample, with progressively fewer active emitters. The acquisition time of each frame was $150$ ms. Only the first and densest frame was used as input for all the studied methods–our CFCNN, R-L deconvolution, and MEF–while the subsequent frames served solely as reference data and were utilized to generate the ground truth, providing a high-resolution benchmark against which the super-resolved outputs of all three methods were compared.
@@ IV.2 Calibration-free convolutional neural network
To train the model, we employed incremental learning, a dynamic approach where newly generated data pairs gradually replace the existing samples during training [44]. This combination of incremental learning and data simulation is particularly beneficial for developing universal networks capable of handling diverse imaging conditions. Throughout 90 incremental iterations, the model processed nearly 500,000 data pairs, with each iteration consisting of 50 epochs of training on mini-batches of 32 samples. The training was performed using the Adam optimizer, which leverages adaptive moment estimation for efficient gradient descent [45]. The initial learning rate was set to $10^{-4}$. Given the high sparsity of the ground truth images, we used an adjusted mean squared error loss function. Firstly, both the ground truth $I$ and the model prediction $\hat{I}$ are convolved with a Gaussian filter $g$ of $\sigma=2$. Then, we compute the squared difference between the filtered images and average it over a batch of $N$ samples, $\frac{1}{N}\sum_{i=1}^{N}\|I_{i}\ast g-\hat{I}_{i}\ast g\|_{2}^{2}$. Similarly, an adjusted mean absolute error metric, $\frac{1}{N}\sum_{i=1}^{N}\|I_{i}\ast g-\hat{I}_{i}\ast g\|_{1}^{2}$, was used to monitor convergence and evaluate the model performance. The presented hyperparameter configuration results from manual optimization of network architecture and dataset ranges to enhance the performance across the device-independent regime. We expect further improvement by using a larger network architecture. However, the computational resources, namely the required operating memory for its training, are presently beyond our technical capacities.
@@ IV.3 Richardson-Lucy algorithm, ThunderSTORM, and DECODE
DECODE is a deep-learning framework for super-resolution fluorescence microscopy that reconstructs emitter positions and intensities by leveraging a convolutional neural network. The utilization of the method is not straightforward, as the user must complete several demanding steps to obtain a model that is not universal and retraining is necessary whenever experimental conditions change. The user trains DECODE for their particular application and setup, which requires a lot of input information. Specifically, a calibration PSF $z$-stack needs to be provided to the external software SMAP, which fits the PSFs, and the fitted parameters are used by DECODE in the training. Therefore, the $z$-stack of a specific PSF must be measured or generated. In addition, the training procedure requires specifying numerous parameters, including emitter density, brightness, camera characteristics, spatial ($x$, $y$, $z$) and intensity ranges, as well as the level of fluorescent background. The training itself typically takes several hours, and the resulting model can only be applied to images with parameters that match those used during training, in all size, shape, and intensity. We trained DECODE separately for our various datasets. In case of the experimental data, we generated a fine-tuned PSF accurately matching the experimentally measured PSFs, and produced the $z$-stack by applying defocus via Zernike polynomials. This was used for the training, along with all specific information about our optical setup. For the synthetic data (Fig. 3 of the main text), we trained DECODE on the precise PSF and two variants of SNR, namely $15$ and $100$. An overall better performance of DECODE was achieved with the latter model, and therefore, we present the corresponding results.
@@
Moreover, each trained DECODE model introduces a systematic shift in the pixel positions of the localized emitters that is consistent across all reconstructions performed by the same model. We corrected for this shift when evaluating the reconstruction quality metrics. Even after applying this correction (which differs between models and is therefore inherently introduced during DECODE’s training) it is evident in the Results chapter that our CFCNN achieves superior performance in all cases.
@@
It is important to note again that all the studied methods were provided only with the single densest fluorescence image of organic molecules. We recognize that both ThunderSTORM and DECODE, as well as our CFCNN, would achieve notably better performance with access to the full sparse image sequence. However, our objective is to compare these methods on dense samples without relying on additional data or imposing specific experimental requirements.
@@ Acknowledgements
AD and RS performed the experiments and numerical simulations. DV developed the deep learning model. AD drafted the manuscript. MJ initiated and supervised the project. All authors contributed in data interpretation and manuscript revisions.

### v3

@@ IV.1 Experiment and sample preparation
The PVA forms an environment where terrylene exhibits very limited photostability, with some of the individual molecules undergoing several blinking cycles before irreversibly photobleaching. We captured a sequence of images from the same location within the sample, with progressively fewer active emitters. The acquisition time of each frame was $150$ ms. Only the first and densest frame was used as input for all the studied methods–our CFCNN, R-L deconvolution, MEF, and DECODE–while the subsequent frames served solely as reference data and were utilized to generate the ground truth, providing a high-resolution benchmark against which the super-resolved outputs of all four methods were compared.
@@ IV.2 Calibration-free convolutional neural network
To train the model, we employed incremental learning, a dynamic approach where newly generated data pairs gradually replace the existing samples during training [45]. This combination of incremental learning and data simulation is particularly beneficial for developing universal networks capable of handling diverse imaging conditions. Throughout 90 incremental iterations, the model processed nearly 500,000 data pairs, with each iteration consisting of 50 epochs of training on mini-batches of 32 samples. The training was performed using the Adam optimizer, which leverages adaptive moment estimation for efficient gradient descent [46]. The initial learning rate was set to $10^{-4}$. Given the high sparsity of the ground truth images, we used an adjusted mean squared error loss function. Firstly, both the ground truth $I$ and the model prediction $\hat{I}$ are convolved with a Gaussian filter $g$ of $\sigma=2$. Then, we compute the squared difference between the filtered images and average it over a batch of $N$ samples, $\frac{1}{N}\sum_{i=1}^{N}\|I_{i}\ast g-\hat{I}_{i}\ast g\|_{2}^{2}$. Similarly, an adjusted mean absolute error metric, $\frac{1}{N}\sum_{i=1}^{N}\|I_{i}\ast g-\hat{I}_{i}\ast g\|_{1}^{2}$, was used to monitor convergence and evaluate the model performance. The presented hyperparameter configuration results from manual optimization of network architecture and dataset ranges to enhance the performance across the device-independent regime. We expect further improvement by using a larger network architecture. However, the computational resources, namely the required operating memory for its training, are presently beyond our technical capacities. The CFCNN model, code, and data are available at the public repository [37].
@@ IV.3 Richardson-Lucy algorithm, ThunderSTORM, and DECODE
DECODE is a deep-learning framework for super-resolution fluorescence microscopy that reconstructs emitter positions and intensities by leveraging a convolutional neural network. The utilization of the method is not straightforward, as the user must complete several demanding steps to obtain a model that is not universal, and retraining is necessary whenever experimental conditions change. The user trains DECODE for their particular application and setup, which requires a lot of input information. Specifically, a calibration PSF $z$-stack needs to be provided to the external software SMAP, which fits the PSFs, and the fitted parameters are used by DECODE in the training. Therefore, the $z$-stack of a specific PSF must be measured or generated. In addition, the training procedure requires specifying numerous parameters, including emitter density, brightness, camera characteristics, spatial ($x$, $y$, $z$) and intensity ranges, as well as the level of fluorescent background. The training itself typically takes several hours, and the resulting model can only be applied to images with parameters that match those used during training, in all size, shape, and intensity. We trained DECODE separately for our various datasets. In case of the experimental data, we generated a fine-tuned PSF accurately matching the experimentally measured PSFs, and produced the $z$-stack by applying defocus via Zernike polynomials. This was used for the training, along with all specific information about our optical setup. For the synthetic data (Fig. 3 of the main text), we trained DECODE on the specific PSF and SNR values separately for each red data point in Fig. 3 (m, n), so that it had the best knowledge of all the parameters possible.
@@
We found that each trained DECODE model introduces a systematic shift in the pixel positions of the localized emitters, which is consistent across all reconstructions performed by the same model. We corrected for this shift when evaluating the reconstruction quality metrics. Even after applying this correction (which differs between models and is therefore inherently introduced during DECODE’s training), it is evident in the Results chapter that our CFCNN achieves superior performance in all cases. Furthermore, the training of the DECODE model is not fully reproducible. When initiated with the same calibration data and the same set of parameters, repeatedly trained DECODE models exhibit varying performance on a test dataset, even though all models pass the DECODE built-in convergence check functionality. In light of this fact, our calibration-free super-resolution approach provides a significant benefit to practitioners in biomedical imaging. Instead of a time-consuming and challenging training process, a user is supplied with an already trained final model, which can be directly applied to a wide variety of microscope setups and samples.
@@
It is important to note again that all the studied methods were provided only with a single densest fluorescence image of organic molecules shown in Fig. 2 (a). We recognize that both ThunderSTORM and DECODE, as well as our CFCNN, would achieve notably better performance with access to the complete sparse image sequence. However, our objective is to compare these methods on dense samples without relying on additional data or imposing specific experimental requirements.
@@ Acknowledgements
AD and RS performed the experiments and numerical simulations. DV developed the deep learning model. AD drafted the manuscript. MJ conceived the idea of calibration-free super-resolution microscopy and supervised the project. All authors contributed to data interpretation and manuscript revisions.